OpenStudy (anonymous):

When a driver presses the brake pedal, his car stops with an acceleration of -8.7 m/s2. How far will the car travel while coming to a complete stop if its initial speed was 10 m/s?

OpenStudy (anonymous):

stopping time, t: v + a t = 10 - 8.7 t = 0 distance = v t + (1/2) a t^2 distance = 10 t -(1/2) 8.7 t^2.